# Dense Video Object Captioning from Disjoint Supervision

### v2

@@ Abstract
Keywords:
vision-and-language object captioning tracking video
@@ 1 Introduction
Figure 1: Overview of the dense video object captioning (Dense VOC) task. Given a video, we predict object trajectories (identities are denoted by colors) and their natural language description. Our task involves object detection, tracking, and captioning. We show a video from the VidSTG [83] validation set.
@@
Powered by gigantic datasets and models, language is becoming the output modality of the most capable artificial intelligence models [57, 10, 1, 46, 35, 42]. Language unifies different tasks with the same output space [49, 11], is more descriptive than discrete class labels [65, 43], and naturally facilitates zero-shot prediction of novel tasks [48, 10]. Inspired by advances in natural language understanding, the vision community has explored language in a number of tasks including image captioning [13], dense image captioning [34], question answering [3], video captioning [45] and representation learning [48]. However, likely due to the scarcity of large-scale, aligned training data, we are not aware of any existing single vision-language model that unifies both fine-grained spatial- (by detecting objects) and temporal- (by reasoning across time in videos) understanding.
@@
In this paper, we propose a new task and model for dense video object captioning (Dense VOC) – the task of generating captions of object trajectories from video (Fig. 1). Dense VOC is therefore a superset of independent tasks commonly studied in vision – namely object detection [24, 41], multi-object tracking [19, 17] and captioning [13, 45], as it requires understanding across space, time and language (Fig. 2).
@@
A prominent challenge for training our model is that datasets with captioned trajectories are scarce. To address this, we design our end-to-end model such that it has separate, but interlinked stages for object proposals, tracking and captioning. Thanks to this design choice, we can train our model without any full annotations for the Dense VOC task by using a mixture of disjoint tasks and datasets that supervise different parts of our model. For example, we can train our object proposal component using image-level object detection labels from COCO [41], and the captioning component from video-level captioning datasets like SMiT [45]. Our disjoint pretraining tasks are complementary, and in combination supervise our entire model. This enables us to perform our Dense VOC task in a zero-shot manner, and we show that we can achieve noteworthy performance despite not having access to any full, captioned object trajectories during training. Furthermore, this pretraining serves as a powerful initialization for finetuning on the full Dense VOC task, where limited annotations are available.
Another challenge in our task is to produce holistic and consistent captions for objects across frames. Note that a baseline of applying a strong, dense image captioning model per-frame, and then linking objects together is poorly suited to this scenario: the captions at each frame are likely to be different due to subtle appearance changes across frames. This motivates our end-to-end trained model, which includes a novel end-to-end tracking algorithm that efficiently aggregates features of the same object across time, enabling the subsequent captioning module to leverage global video features to produce coherent captions.
To evaluate our model, we develop a new metric that jointly measures performance on captioning, detection and tracking by extending HOTA [44], the most popular metric for multi-object tracking. Although we are the first to our knowledge to study Dense VOC, we can still repurpose existing video grounding datasets for evaluation and domain-specific finetuning. We use VidSTG [83] and VLN [60], originally designed for spatiotemporal sentence grounding: Instead of finding an object tube given a sentence query (grounding), we predict object trajectories directly and use the sentence queries as the ground truth captions.
@@
Figure 2: Overview of Dense VOC. Our problem involves understanding across space, time and language, and thus encompasses other vision tasks, which typically consider one or two of these axes. We also show how these subtasks are complementary, and how pretraining on them enables zero-shot generalization to Dense VOC.
In addition, we show that our generative model trained for Dense VOC can perform grounding by simply selecting the bounding boxes with the maximum likelihood of producing the query sentence.
@@
Experiments show that our end-to-end trained Dense VOC model outperforms baselines consisting of strong, per-task models by a substantial margin, producing more accurate and inherently temporally consistent captions. Moreover, we achieve significant improvements from our disjoint, multi-dataset training. Furthermore, by applying our generative captioning model to the discriminative grounding task, we are able to outperform dedicated spatial grounding models on both VidSTG and VLN. In summary, we propose the following contributions:
@@
    We further show how our models generalize to downstream grounding tasks, achieving state-of-the-art results on two datasets, without explicitly being trained on grounding tasks.
@@ 2 Related Work
Image captioning [13, 2, 69, 52] describes the content of an image with language. State-of-the-art methods map the input image to output text by using multi-modal models [30, 20, 38, 80, 35, 78] pretrained on large datasets [55, 48]. For example, GIT [61] simple forwards vision tokens from a ViT encoder [22] to an auto-regressive language decoder [59, 21]. Similar ideas apply to video captioning [68, 85, 45], by concatenating [61] or pooling [71] features from each frame, before feeding them to the auto-regressive text decoder. Our work builds on existing captioning architectures [61], and extends them to object trajectory captioning using our end-to-end model and weak supervision [45, 34, 41].
@@
Dense object captioning in contrast, detects objects in an image and describes them with text [32, 37, 54, 65]. It was popularized by the Visual Genome [34] dataset, which contains full annotations for the task. Early work, DenseCap [32] used a one-stage detector [50] followed by an LSTM text decoder [29] on dense feature maps. Most recently, GRiT [65] built upon the state-of-the-art image captioning architecture of GIT [61], and generated object captions, also with a transformer decoder [59], from RoI-pooled [27] image features. Our model advances architectures like GRiT to videos and incorporates end-to-end tracking of object trajectories too. We also note that dense video captioning in the literature refers to the task of localizing and captioning multiple events temporally in videos [33, 85, 62, 73]. Our task, in contrast, involves tracking and captioning objects in a video, and therefore requires spatial localization, which is why we name our task “dense video object captioning”.
@@
Multi-object tracking aims to detect objects and track them with a consistent identity label. The predominant approach is tracking-after-detection [7, 82, 23], i.e. first running detectors on each frame and then using a separate tracker to link them. While this works well for existing benchmarks with only a few classes [19, 25, 74], it is more challenging in our case: we need tracks before captioning to have a single, consistent textual output for the whole trajectory. Thus, our work follows end-to-end multi-object tracking [14, 36, 64, 89]. We adopt a global tracker GTR [89], which casts tracking as pairwise association among all objects within a video. Whilst GTR applies a sliding-window-based identity association algorithm during inference as a post-processing step, we design an efficient algorithm to perform this process end-to-end. This is necessary for our task, since our trajectory features are used by a subsequent captioning module which is trained jointly. We are not aware of prior work which efficiently assigns object identities and corresponding features to tracks, and trains end-to-end through this process. Finally, we note that video object tracking and segmentation [76, 75, 77, 16, 15] focuses on following a single object which is given in the first frame [47, 70]. This is therefore a different setting from our task of detecting, tracking and captioning multiple objects.
@@
Video object grounding [83, 60] finds a spatio-temporal tube given a video and query sentence as inputs. Existing, discriminative methods [83, 72, 31, 56] co-embed visual and text inputs, and use the sentence feature to find the corresponding object. In contrast, we use our generative language model for this task by selecting the object with the highest likelihood of producing the query. To our knowledge, we are the first work to explore the alternate paradigm of generative models for this task. Finally, we note that these tasks are also related to video-referring segmentation [6, 66, 79] which grounds textual queries to segmentation masks. Segmentation, however, is not the focus of our work.
@@ 3 Method
Figure 3: Overview of our model. Our end-to-end model has three modules: First it produces object proposals per-frame using a class-agnostic detector (left, trained with detection loss, $L_{object}$). These object proposals are then passed to an end-to-end tracking module that groups objects into trajectories (middle, trained with association loss, $L_{assoc}$). The identities produced by the tracking module are used to aggregate features which are then fed to a language decoder to produce the final caption (right, trained with caption loss $L_{caption}$). Our full model can be trained end-to-end with Dense VOC supervision when available, or be trained with partial supervision on different and disjoint datasets to provide zero-shot capabilities, and initialization for finetuning.
As shown in Fig. 3, our end-to-end trainable model consists of interlinked heads for object proposal, tracking and captioning the resulting trajectories. Before introducing our novel components, we review prior techniques for captioning and dense object captioning in images [65, 61] which our model is built on.
@@ 3.1 Background
Image captioning maps an input image, $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, to a caption $c=(y_{1},y_{2},\ldots,y_{n_{t}})$ which is a sequence of up to $n_{t}$ text tokens from a given vocabulary. The minimal set of components is an image encoder, followed by a text decoder [59]. The encoder maps the input image $\mathbf{I}$, to a feature representation $\mathbf{f}\in\mathbb{R}^{n_{v}\times d}$ consisting of $n_{v}$ tokens with dimensionality $d$. The subsequent text decoder is auto-regressive [59, 26] – it predicts the next text token, $y_{i}$, as a function of both the image features, $\mathbf{f}$, and previously generated text tokens, $\mathbf{y}_{0:i-1}$, denoted by $y_{i}=\text{Decode}(\mathbf{f},\mathbf{y}_{0:i-1})$. Note that the first step of decoding begins with $y_{0}=\texttt{BOS}$, a special beginning-of-sentence token, and the caption ends when the end-of-sentence token, EOS, is output by the model. This simple image captioning model has been demonstrated to be effective and scalable by GIT [61], achieving state-of-the-art results across a number of captioning datasets.
GRiT [65] extends the approach further to dense object captioning of images: Here, the authors use an object proposal network [88] to produce a set of $K$ class-agnostic bounding boxes, $b_{1},b_{2},\ldots,b_{K}$. Features corresponding to each of these objects are obtained using RoIAlign [27], resulting in a localized feature, $f_{k}\in\mathbb{R}^{r\times r\times d}$ where $r=7$ is the output resolution of RoIAlign. Each of these grid features is flattened into $f_{k}\in\mathbb{R}^{r^{2}\times d}$ and decoded independently by the text decoder, as done in GIT [61]. Therefore, the loss used to train a GRiT model consists of $L=L_{object}+L_{caption}$ where $L_{caption}$ is a cross-entropy loss over all text tokens in the vocabulary, and $L_{object}$ consists of bounding box regression and objectness terms, as standard in object detection literature [88, 51, 40].
We now describe how we extend object captioning to videos by tracking object proposals over time (Sec. References) and aggregating trajectory features and captioning them (Sec. References) in an end-to-end fashion. Section References explains how we train our model, whilst Sec. References describes how we apply our model directly to video object grounding tasks.
@@ 3.2 End-to-end tracking
From these object features, $\mathbf{F}$, we predict a global association matrix, $\mathbf{A}\in\mathbb{R}^{M\times M}$, where $\mathbf{A}_{ij}=1$ if the objects denoted by the $i^{th}$ row and $j^{th}$ column are from the same trajectory (Fig. 3 middle). Otherwise, $\mathbf{A}_{ij}=0$ meaning that these objects are from different trajectories, or at least one of them is background.
We use a transformer module, H, with two self-attention layers, similar to GTR [89], to predict the association matrix $\mathbf{A}=\sigma(\text{H}(\mathbf{F})$), where $\sigma$ is the sigmoid activation. Given the object trajectory annotations, we construct the ground truth association matrix $\mathbf{\bar{A}}$ for $\mathbf{A}$, where $\mathbf{\bar{A}}_{ij}=1$ if and only if row $i$ and column $j$ of $\mathbf{A}$ are matched to the same ground truth trajectory using an Intersection over Union (IoU) criteria of 0.5. The training loss $L_{assoc}$ for this module is then a binary cross entropy between $\mathbf{A}$ and $\mathbf{\bar{A}}$, $L_{assoc}=\frac{1}{M}\sum_{ij}\text{BCE}(A_{ij},\bar{A}_{ij})$.
Input :  Association Matrix $\bf{A}\in\mathbb{R}^{TK\times TK}$, where $T$ denotes the number of frames, and $K$ the number of objects per frame.
@@
$A\leftarrow$ preprocess($A$) // Preprocess $A$ to ensure object pairs
in the same frame have a score of 00.
@@
       track_len $\leftarrow$ $\hat{A}$.sum(axis=1) // Compute the number of objects in each merge.
@@
       $\delta\leftarrow\delta+$ id_count * $\hat{A}_{i}$ // Assign the current track a new ID
       using $\hat{A}_{i}$ as the binary mask.
@@
Algorithm 1 Identity assignment from association matrix. This greedy algorithm can be implemented efficiently on accelerators, allowing us to train end-to-end with it.
After constructing our association matrix, $\mathbf{A}$, we need to aggregate object-level features according to identities $\delta=[\delta_{k}^{t}]_{t=1,k=1}^{T,K_{t}}$, to generate trajectory-level captions for the next captioning stage. Here, $\delta_{k}^{t}$ denotes the identity of the $k$-th object proposal in the $t$-th frame. We design a greedy grouping algorithm (Alg. 1) operating on $\mathbf{A}$ to obtain $\delta$. Concretely, we greedily extract the longest trajectory from untracked objects, until there are no possible associations left (indicated by the association score being above a threshold $\theta$). This guarantees that each trajectory has at most one object in each frame. This algorithm can be implemented efficiently on accelerators, allowing us to backpropagate through it.
As aforementioned, prior trackers [82, 87, 86] do not explicitly perform identity assignment within the model, but rather as a post-processing step since tracking is the final output for such methods. Our work efficiently assigns object identities to tracks in an end-to-end trainable network, which enables us to perform joint trajectory-level captioning training as described next.

### v3

@@ 1 Introduction
Powered by gigantic datasets and models, language is becoming the output modality of the most capable artificial intelligence models (Team et al., 2023; Alayrac et al., 2022; Ouyang et al., 2022; Li et al., 2023; Liu et al., 2023; Tong et al., 2024; Li et al., 2024a). Language unifies different tasks with the same output space (Raffel et al., 2020; Chen et al., 2023a), is more descriptive than discrete class labels (Wu et al., 2022a; Long et al., 2023), and naturally facilitates zero-shot prediction of novel tasks (Radford et al., 2021; Brown et al., 2020). Inspired by advances in natural language understanding, the vision community has explored language in a number of tasks including image captioning (Chen et al., 2015), dense image captioning (Krishna et al., 2017b), question answering (Antol et al., 2015), video captioning (Monfort et al., 2021) and representation learning (Radford et al., 2021). However, likely due to the scarcity of large-scale, aligned training data, we are not aware of any existing single vision-language model that unifies both fine-grained spatial- (by detecting objects) and temporal- (by reasoning across time in videos) understanding.
@@
In this paper, we propose a new task and model for dense video object captioning (Dense VOC) – the task of generating captions of trajectories of all objects from video (Fig. 1). Dense VOC requires understanding across space, time, and language (Fig. 2), and is therefore a superset of existing vision tasks, namely object detection (Everingham et al., 2015; Lin et al., 2014), multi-object tracking (Dendorfer et al., 2021; Dave et al., 2020) and captioning (Chen et al., 2015).
Figure 1: Overview of the dense video object captioning (Dense VOC) task. Given a video, we predict object trajectories (identities denoted by colors) and their natural language description. We show a video from the VidSTG (Zhang et al., 2020) validation set.
@@
A prominent challenge for training our model is that datasets with captioned trajectories are scarce. However, annotations for each sub-task, or even each combination of the sub-tasks, are abundant. For example, we can train our object proposal component using image-level object detection labels from COCO (Lin et al., 2014), and the captioning component from video-level captioning datasets like SMiT (Monfort et al., 2021). These disjoint training tasks are complementary, and in combination supervise our entire model. This enables us to perform our Dense VOC task in a zero-shot manner, and we show that we can achieve noteworthy performance despite not having access to any full, captioned object trajectories during training. Furthermore, this pretraining serves as a powerful initialization for finetuning on the full Dense VOC task, where limited annotations are available.
@@
Figure 2: Overview of Dense VOC. Our problem involves understanding across space, time, and language, and thus encompasses other vision tasks, which typically consider one or two of these axes. We show these subtasks are complementary, and pretraining on them enables zero-shot generalization to Dense VOC.
Another challenge in our task is to produce holistic and consistent captions for objects across frames. Note that a baseline of applying a strong, dense image captioning model per-frame, and then linking objects together is poorly suited to this scenario: the captions at each frame are likely to be different due to subtle appearance changes across frames. This motivates our end-to-end trained model, which includes a novel end-to-end tracking algorithm that aggregates features of the same object across time, enabling the subsequent captioner to leverage global features to produce coherent captions.
Although we are the first to our knowledge to study Dense VOC, we can still repurpose existing video grounding datasets for evaluation and domain-specific finetuning. We use VidSTG (Zhang et al., 2020) and VLN (Voigtlaender et al., 2023), originally designed for spatiotemporal sentence grounding: Instead of finding an object tube given a sentence query (grounding), we predict object trajectories directly and use the sentence queries as the ground truth captions. In addition, we show that our generative model trained for Dense VOC can perform grounding by simply selecting the bounding boxes with the maximum likelihood of producing the query sentence. We also develop a new metric that jointly measures captioning, detection and tracking accuracy by extending HOTA (Luiten et al., 2021), the most popular metric for multi-object tracking.
@@
Experiments show that our end-to-end trained Dense VOC model outperforms baselines consisting of strong, per-task models by a substantial margin, producing more accurate and inherently temporally consistent captions. Moreover, we achieve significant improvements from our disjoint, multi-dataset training. We additionally show how we can readily apply our model to related domain-specific datasets: by finetuning our model on a recent person tracking and captioning dataset, BenSMOT (Li et al., 2024b), we outperform prior work by $18.2$ points. Furthermore, by applying our generative captioning model to the discriminative grounding task, we are able to outperform dedicated spatial grounding models on both VidSTG and VLN. In summary, we propose the following contributions:
@@
    We further show how our models generalize to downstream video grounding tasks, achieving state-of-the-art results on two datasets, without explicitly being trained for grounding.
  5. 5.
    Moreover, we significantly improves the state-of-the-art on the BenSMOT dataset Li et al. (2024b) for Semantic Multi-Object Tracking.
@@ 2 Related Work
Image captioning (Chen et al., 2015; Anderson et al., 2018; Xu et al., 2015; Rennie et al., 2017) describes the content of an image with language. State-of-the-art methods map the input image to output text by using multi-modal models (Jiang et al., 2020; Desai & Johnson, 2021; Li et al., 2020; Zhang et al., 2021a; Li et al., 2023; Yu et al., 2022) pretrained on large datasets (Sharma et al., 2018; Radford et al., 2021). For example, GIT (Wang et al., 2022) simple forwards vision tokens from a ViT encoder (Dosovitskiy et al., 2021) to an auto-regressive language decoder (Vaswani et al., 2017; Devlin et al., 2019). Similar ideas apply to video captioning (Xu et al., 2016; Zhou et al., 2018; Monfort et al., 2021), by concatenating (Wang et al., 2022) or pooling (Yan et al., 2022) features from each frame, before feeding them to an auto-regressive text decoder. Our work builds on existing captioning architectures (Wang et al., 2022), and extends them to trajectory captioning using our end-to-end model and weak supervision (Monfort et al., 2021; Krishna et al., 2017b; Lin et al., 2014).
@@
Dense object captioning in contrast, detects objects in an image and describes them with text (Johnson et al., 2016; Li et al., 2019; Shao et al., 2022; Wu et al., 2022a). It was popularized by the Visual Genome (Krishna et al., 2017b) dataset, which contains full annotations for the task. Early work, DenseCap (Johnson et al., 2016) used a one-stage detector (Redmon et al., 2016) followed by an LSTM text decoder (Hochreiter & Schmidhuber, 1997) on dense feature maps. Most recently, GRiT (Wu et al., 2022a) built upon the state-of-the-art image captioning architecture of GIT (Wang et al., 2022), and generated object captions, also with a transformer decoder (Vaswani et al., 2017), from RoI-pooled (He et al., 2017) image features. Our model advances architectures like GRiT to videos and incorporates end-to-end tracking. We also note that dense video captioning in the literature refers to the task of localizing and captioning multiple events temporally in videos (Krishna et al., 2017a; Zhou et al., 2018; Wang et al., 2021a; Yang et al., 2023a). Our task, in contrast, involves tracking and captioning objects in a video, and therefore requires spatial localization, which is why we name our task “dense video object captioning”.
@@
Multi-object tracking detects objects and track them with a consistent identity label. The predominant approach is tracking-after-detection (Bewley et al., 2016; Zhang et al., 2021b; Du et al., 2021), i.e. first running detectors on each frame and then using a separate tracker to link them. While this works well for existing benchmarks with only a few classes (Dendorfer et al., 2021; Geiger et al., 2012; Yang et al., 2019), it is more challenging in our case: we need tracks before captioning to have a single, consistent textual output for the whole trajectory. Thus, our work follows end-to-end multi-object tracking (Cheng et al., 2022; Li et al., 2022a; Wang et al., 2021c; Zhou et al., 2022b). We adopt a global tracker GTR (Zhou et al., 2022b), which casts tracking as pairwise association among all objects within a video. Whilst GTR applies a sliding-window-based identity association algorithm during inference as a post-processing step, we design an efficient algorithm to perform this process end-to-end. This is necessary for our task, since our trajectory features are used by a subsequent captioning module which is trained jointly. We are not aware of prior work which efficiently assigns object identities and corresponding features to tracks, and trains end-to-end through this process. Finally, note that video object tracking and segmentation (Yang et al., 2021; 2023b; Yang & Yang, 2022; Cheng & Schwing, 2022; Cheng et al., 2024) focuses on following only a single object which is given in the first frame (Perazzi et al., 2016; Xu et al., 2018). This is therefore a different setting from our task of detecting, tracking and captioning multiple objects.
@@
Video object grounding (Zhang et al., 2020; Voigtlaender et al., 2023) finds a spatio-temporal tube given a video and query sentence as inputs. Existing, discriminative methods (Zhang et al., 2020; Yang et al., 2022; Jin et al., 2022; Su et al., 2021) co-embed visual and text inputs, and use the sentence feature to find the corresponding object. In contrast, we use our generative language model for this task by selecting the object with the highest likelihood of producing the query. To our knowledge, we are the first work to explore the alternate paradigm of generative models for this task. Finally, we note that these tasks are also related to video-referring segmentation (Bellver et al., 2020; Wu et al., 2022b; Yu et al., 2016) which grounds textual queries to segmentation masks. Segmentation, however, is not the focus of our work.
Concurrent to our work, BeyondMOT (Li et al., 2024b) proposes an video object tracking and captioning benchmark and model. We highlight two differences: 1. Li et al. (2024b) uses a frame-by-frame tracker similar to our baselines (Tab. 2), and we propose a novel end-to-end tracker. 2. Our work aims to track and caption all objects in the video, while Li et al. (2024b) handles only persons. As a result, our task is much more challenging, and we show our model yields superior performance on their benchmark. OW-VISCap (Choudhuri et al., 2024) on the other hand augments a video segmentation model, Cheng et al. (2022), with a language model (OPT with 2.7 billion parameters (Zhang et al., 2022a)) head for video segmentation and captioning. In contrast, our model is trained flexibly using our disjoint pretraining, which enables us to achieve better detection and tracking performance whilst still using a substantially smaller model.
@@ 3 Method
Figure 3: Overview of our model. Our end-to-end model has three modules: First it produces object proposals per-frame using a class-agnostic detector (left, trained with detection loss, $L_{object}$). These object proposals are then passed to an end-to-end tracking module that groups objects into trajectories (middle, trained with association loss, $L_{assoc}$). The identities produced by the tracking module are used to aggregate features which are then fed to a language decoder to produce the final caption (right, trained with caption loss $L_{caption}$). Our model can be trained end-to-end with partial supervision on different and disjoint datasets to provide zero-shot Dense VOC capabilities.
As shown in Fig. 3, our end-to-end model consists of interlinked heads for object proposal, tracking and captioning the resulting trajectories. Before introducing our novel components, we review prior techniques for captioning and dense object captioning in images (Wu et al., 2022a; Wang et al., 2022).
@@ 3.1 Background
Image captioning maps an input image, $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, to a caption $c=(y_{1},y_{2},\ldots,y_{n_{t}})$ which is a sequence of up to $n_{t}$ text tokens from a given vocabulary. The minimal set of components is an image encoder, followed by a text decoder (Vaswani et al., 2017). The encoder maps the input image $\mathbf{I}$, to a feature representation $\mathbf{f}\in\mathbb{R}^{n_{v}\times d}$ consisting of $n_{v}$ tokens with dimensionality $d$. The subsequent text decoder is auto-regressive (Graves, 2013) – it predicts the next text token, $y_{i}$, as a function of both the image features, $\mathbf{f}$, and previously generated text tokens, $\mathbf{y}_{0:i-1}$, denoted by $y_{i}=\text{Decode}(\mathbf{f},\mathbf{y}_{0:i-1})$. Note that the first step of decoding begins with $y_{0}=\texttt{BOS}$, a special beginning-of-sentence token, and the caption ends when the end-of-sentence token, EOS, is output by the model. This simple image captioning model has been demonstrated to be effective and scalable by GIT (Wang et al., 2022), achieving state-of-the-art results across a number of captioning datasets.
GRiT (Wu et al., 2022a) extends the approach further to dense object captioning of images: Here, the authors use an object proposal network (Zhou et al., 2019) to produce a set of $K$ class-agnostic bounding boxes, $b_{1},b_{2},\ldots,b_{K}$. Features corresponding to each of these objects are obtained using RoIAlign (He et al., 2017), resulting in a localized feature, $f_{k}\in\mathbb{R}^{r\times r\times d}$ where $r=7$ is the output resolution of RoIAlign. Each of these grid features is flattened into $f_{k}\in\mathbb{R}^{r^{2}\times d}$ and decoded independently by the text decoder, as done in GIT. Therefore, the loss used to train a GRiT model consists of $L=L_{object}+L_{caption}$ where $L_{caption}$ is a cross-entropy loss over all text tokens in the vocabulary, and $L_{object}$ consists of bounding box regression and objectness terms, as standard in object detection literature (Zhou et al., 2019; Ren et al., 2015; Lin et al., 2017).
We now describe how we extend object captioning to videos by tracking object proposals over time (Sec. 3.2) and aggregating trajectory features and captioning them (Sec. 3.3) in an end-to-end fashion. Section 3.4 explains how we train our model, whilst Sec. 3.5 describes how we apply our model directly to video object grounding tasks.
@@ 3.2 End-to-end tracking
Input :  Association Matrix $\bf{A}\in\mathbb{R}^{TK\times TK}$ // $T$: num. frames. $K$: num. objects per frame.
@@
$A\leftarrow$ preprocess($A$) // Preprocess $A$ to ensure object pairs in the same frame have a score of 00.
@@
       track_len $\leftarrow$ $\hat{A}$.sum(axis=1) // Number of objects in each merge.
@@
       $\delta\leftarrow\delta+$ id_count * $\hat{A}_{i}$ // Assign the current track a new ID using $\hat{A}_{i}$.
@@
Algorithm 1 Identity assignment from association matrix. This greedy algorithm can be implemented efficiently on accelerators, enabling end-to-end training.
From these object features, $\mathbf{F}$, we predict a global association matrix, $\mathbf{A}\in\mathbb{R}^{M\times M}$, where $\mathbf{A}_{ij}=1$ if the objects denoted by the $i^{th}$ row and $j^{th}$ column are from the same trajectory (Fig. 3 middle). Otherwise, $\mathbf{A}_{ij}=0$ means that they are from different trajectories, or one of them is the background.
We use a transformer module, H, with two self-attention layers, similar to  Zhou et al. (2022b), to predict the association matrix $\mathbf{A}=\sigma(\text{H}(\mathbf{F})$), where $\sigma$ is the sigmoid activation. Given the object trajectory annotations, we construct the ground truth association matrix $\mathbf{\bar{A}}$ for $\mathbf{A}$, where $\mathbf{\bar{A}}_{ij}=1$ if and only if row $i$ and column $j$ of $\mathbf{A}$ are matched to the same ground truth trajectory using an Intersection over Union (IoU) criteria of 0.5. The training loss $L_{assoc}$ for this module is then a binary cross entropy between $\mathbf{A}$ and $\mathbf{\bar{A}}$, $L_{assoc}=\frac{1}{M}\sum_{ij}\text{BCE}(A_{ij},\bar{A}_{ij})$.
After constructing our association matrix, $\mathbf{A}$, we need to aggregate object-level features according to identities $\delta=[\delta_{k}^{t}]_{t=1,k=1}^{T,K_{t}}$, to generate trajectory-level captions for the next captioning stage. Here, $\delta_{k}^{t}$ denotes the identity of the $k$-th object proposal in the $t$-th frame. We design a greedy grouping algorithm (Alg. 1) operating on $\mathbf{A}$ to obtain $\delta$. Concretely, we greedily extract the longest trajectory from untracked objects, until there are no possible associations left (indicated by the association score being above a threshold $\theta$). This guarantees each trajectory has at most one object in each frame. This algorithm can be implemented efficiently on accelerators, allowing us to backpropagate through it.
As aforementioned, prior trackers (Zhang et al., 2021b; Zhou et al., 2020; 2022a) do not explicitly perform identity assignment within the model, but rather as a post-processing step since tracking is the final output for such methods. Our work efficiently assigns object identities to tracks in an end-to-end trainable network, which enables us to perform joint trajectory-level captioning training as described next.